OpenStudy (awesomedude2078):

Use this data in the problem below. Follow the steps carefully. Round to the nearest tenth. Lot 1: Week 1: 600 Week 2: 300 Week 3: 500 Week 4: 900 Step 1. Jim enters the data and calculates the average or mean. _____________ Step 2. Jim calculates the deviation from the mean by subtracting the mean from each value. Step 3. Jim squares each deviation to remove negative signs. Step 4. Jim sums the squares of each deviation and divides by the count for the variance. _____________ Step 5. Jim takes the square root of the variance to find the standard deviation. _____________

OpenStudy (otherworldly):

do u know what mean is

OpenStudy (otherworldly):

For step 1 to get the mean/average u have to add all the numbers together than divide by the number of numbers u added so if u have 4 numbers divide the sum by it

OpenStudy (awesomedude2078):

600+300+500+900=2300/4=575 is the average or mean

OpenStudy (otherworldly):

great for step two is there like a space for the answer or is it just saying that u don't need to answer anything?

OpenStudy (awesomedude2078):

only answers in the blanks

OpenStudy (otherworldly):

k

OpenStudy (awesomedude2078):

25^2+(-275^2)+(-75^2)+325^2= right

OpenStudy (otherworldly):

yd u put to the power of two

OpenStudy (awesomedude2078):

he squares them to get rid of the negatives

OpenStudy (otherworldly):

squaring has nothing to do with the power of two

OpenStudy (otherworldly):

\[\sqrt{25}\]

OpenStudy (otherworldly):

that would be 5

OpenStudy (otherworldly):

\[\sqrt{?} \]

OpenStudy (otherworldly):

this is the symbol of squaring

OpenStudy (otherworldly):

Do u get it?

OpenStudy (otherworldly):

pellet nvm wow I'm supposed to be helping u and I'm confusing u and me damn i really do suk at explaining

OpenStudy (otherworldly):

sh!t*

OpenStudy (otherworldly):

srry i don't think i should explain this I'm kinda failing